Solve the simultaneous equations


2x^2 = 4y and logxy = 3/2 


1
Expert's answer
2021-08-02T16:41:11-0400

Explanations & Calculations


"\qquad\qquad
\begin{aligned}
\small 2^{x^2}&=\small 4^y\\
\small 2^{x^2}&=\small (2^2)^y= 2^{2y}\\\\
\text{By equaling the indeces we get}\\\\
\small x^2&=\small 2y\cdots\cdots\cdots\cdots(1)\\\\

\small log_x\,y&=\small \frac{3}{2}\\
\small y&=\small x^{\frac{3}{2}}\cdots\cdots\cdots\cdots(2)\\\\

\small\text{By (1)=(2),}\\\\

\small x^2 &=\small 2x^{\frac{3}{2}}\\
\small x^2-2x^{\frac{3}{2}}&=\small 0\\
x^{\frac{3}{2}}(x^{\frac{1}{2}}-2)&=\small 0\\
\small x^{\frac{3}{2}} = 0\qquad &or\qquad x^{\frac{1}{2}}-2=0\\
\small x= 0\qquad &or\qquad x =2^2=4 \\\\

\small \text{accordingly the y values by (2),}\\\\

\small y = 0\qquad & or \qquad y= 8\\\\

\small \therefore (x,y) = (0,0)\qquad &or\qquad (4,8)

\end{aligned}"


  • All x & y values satisfy the above (1) & (2) equations hence form the results.